100 mL of 0.1M H2SO4 and 100 mL of 0.1M
 
KOH are mixed together and diluted to
exactly two litres. Normality of unreacted
acid in the mixture is

Grade:12

2 Answers

Pooja
4 years ago
For H2SO4
 H2SO4= 2H+SO4-2
Concentration of Hions is
100× 0.1 ×2=20millimoles
 
For KOH
KOH=K+ OH-
 
Pooja
4 years ago
Cont. (Sorry I accidentally submitted incomplete answer)
Concentrationof KOH is 100 × 0.1=10millimoles
 
Now after reaction the millimole left are 20-10 = 10millimoles 
Normality of remaining acid in 2 litre volume is
10/2 = 5millinormal